OpenStudy (anonymous):

Can someone explain me this word problem the length of a rectangle is 3 more than twice the width, w. Write an equation giving the area, A, as a function of the width.

OpenStudy (shamim):

Let the length=l

OpenStudy (shamim):

So l=2w+3 Right?

OpenStudy (shamim):

U know Area of a rectangle=length*weidth

OpenStudy (shamim):

Area=l*w

OpenStudy (shamim):

Put the value of l

OpenStudy (shubhipapa):

The equation of area=(2w+3)w =2w^2+3w

OpenStudy (think2raj):

area (A)=3*1=3

Directrix (directrix):

The area as a function of w would be written as: A(w) = w*(3 + 2*w)